# Practical Aspects on Solving Differential Equations Using Deep Learning: A Primer

### v3

@@ 3.1 Feed-forward neural network
Computational units or neurons lie within each layer, performing affine transformations. We can write that in a vector form as ${\bf y}_{l}=f_{l}({\bf z})=g_{l}({\bf W}_{l}\cdot{\bf z}+{\bf b}_{l})$, where ${\bf W}_{l}$ contains the weights that connect neurons from layer $l-1$ with neurons in layer $l$ and ${\bf b}_{l}$ are the biases of the neurons at layer $l$, and number of layers $l$ runs $l=1,\ldots,k$. The function $g(x)$ is called the activation function and can be (i) a Sigmoid $g(x)=\frac{1}{1+\exp(-x)}$, (ii) a hyperbolic tangent $g(x)=\tanh(x)$, a (iii) rectified linear unit (ReLU) $g(x)=\max\{x,0\}$, any custom function, or any of the functions that the reader can find in Pytorch documentation page [^3] https://pytorch.org/docs/stable/nn.html##non-linear-activations-weighted-sum-nonlinearity.
@@
Like all the modern deep learning libraries, Pytorch offers all the necessary tools to build feed-forward neural networks and define affine transformations easily. The reader can find a complete list on the PyTorch documentation page [^4] https://pytorch.org/docs/stable/nn.html#linear-layers. Moreover, code listing 1 shows a basic implementation of a Pytorch Multi-Layer Perceptron neural network (a feed-forward neural network) with linear layers and $\tanh$ activation functions. We must define the layers and the activation functions used in the constructor __init__ method. The same method receives as arguments the input and output dimensions, which are the number of independent and dependent variables in our case, respectively. The argument input_dim defines the number of input units (independent variables such as $t$ and ${\bf x}$), and on the other hand, the argument output_dim determines the number of units in the output layer (e.g., $y$). Moreover, it receives the number of layers (num_layers) and the number of units in each hidden layer (hidden_size). In this implementation, we assume that all hidden layers will have the same number of units. Keep in mind that if one has to use a neural network with two hidden layers, they have to set the number of layers to zero, for two hidden layers, the number has to be set to one, and so on. In line $11$ of code listing 1, we define the input layer; in lines 1414141415151515, the hidden layers; and in line $18$, the output layer. The class Linear of Pytorch is an implementation of an affine transformation (${\bf y}={\bf W}\cdot{\bf x}+{\bf b}$). Notice that in line $14$, we use a ModuleList to instantiate the hidden layers. We do so because the number of hidden layers is not known a priori; instead, it is passed as an argument to the constructor when we instantiate the MLP class. The forward method takes the input tensor as an argument (it might accept more arguments, as we will see later) and applies all the transformations and activation functions we have already defined. Essentially, this method runs the forward pass of our neural network, which is the part of the code where we determine the structure of our neural network. Keep in mind that when we use the batch normalization (batch_norm = True), we can eliminate the bias term from the previous Layer by setting bias=False in the constructor of the Layer class in Pytorch.
1class MLP(nn.Module):
@@
11 self.fc_in = nn.Linear(input_dim, hidden_size)
@@
14 self.layers = nn.ModuleList([nn.Linear(hidden_size, hidden_size)
@@
18 self.fc_out = nn.Linear(hidden_size, output_dim)
@@
25 self.bn = nn.BatchNorm1d(hidden_size)
@@ 3.2 DGM neural network
1 class DGMLayer(nn.Module):
@@
5 self.Z_wg = nn.Linear(hidden_size, hidden_size)
6 self.Z_ug = nn.Linear(input_dim, hidden_size, bias=False)
@@
8 self.G_wz = nn.Linear(hidden_size, hidden_size)
9 self.G_uz = nn.Linear(input_dim, hidden_size, bias=False)
@@
11 self.R_wr = nn.Linear(hidden_size, hidden_size)
12 self.R_ur = nn.Linear(input_dim, hidden_size, bias=False)
@@
14 self.H_wh = nn.Linear(hidden_size, hidden_size)
15 self.H_uh = nn.Linear(input_dim, hidden_size, bias=False)
@@
18 self.sigma = nn.Tanh()
@@ 3.3 Feed-forward ResNet
An example of implementing a linear ResNet block in Pytorch is given in code listing 3. The method __init__ takes the number of units in the input and output layers as arguments, respectively. The argument downsample defines a transformation (i.e., ${\bf W}_{s}$) at the skip connections that we will apply when the input’s ${\bf x}$ dimensions mismatches the size of the output of the residual block (see Figure 2 C). Finally, the argument num_features indicates if the input tensor will contain an extra dimension besides the batch size and the independent variables. Notice that the linear layers here are defined using the Pytorch Sequential container in lines 1010101014141414. Within each container, we instantiate a Pytorch linear layer followed by a batch normalization layer (see Section 3.7). In the forward method in lines 2323232333333333, we apply two linear layers (see lines $20$ and $21$ in code listing 3) and linear rectifications on their outputs. Then we check if downsample is not None (see lines $23$ and $24$), and if it’s not, then we apply the downsample transformation to the input ${\bf x}$, otherwise our residual remains the original input ${\bf x}$. Once we have the residual values, we add them to the output, apply linear rectification, and return the output.
1class ResidualBlock(nn.Module):
@@
15 self.fc1 = nn.Sequential(nn.Linear(input_dim, output_dim, bias=False),
16 nn.BatchNorm1d(self.num_features))
@@
18 self.fc2 = nn.Sequential(nn.Linear(output_dim, output_dim, bias=False),
19 nn.BatchNorm1d(self.num_features))
@@
21 self.relu = nn.ReLU()
@@ 3.5 Backpropagation, Optimizers and Automatic Differentiation
Algorithm 1 summarizes the learning steps, and code listing 4 shows the backbone of training a neural network in Pytorch. In general, we first instantiate our neural network class (line $2$ in code listing 4), then we specify the optimizer (in this case, at line $4$ we use the Adam optimizer [40]), and the criterion, which is the loss function we are interested in minimizing. In this work, we define custom loss functions per equation; we will see in Section 4.1 how to implement a custom loss function. Moreover, we might want to initialize a learning rate scheduler, which will change (usually decrease) the learning rate based on some linear, exponential, sinusoidal, or other relation [^6] The reader can find a list of learning rate schedulers already implemented in PyTorch here: https://pytorch.org/docs/stable/optim.html#how-to-adjust-learning-rate. Then, for each iteration (line $7$), we sample our independent variables (e.g., time $t$, space $x$, or both) based on predefined distributions. Once we get the samples, we zero the grads (line $10$), then we approximate the solution yhat using the neural network and the independent variables. Finally, we pass the approximation to the loss function, do a backward step (lines $13$-$18$), and finally minimize it (line $21$).
@@ 3.7 Batch Normalization
Let =$\mathcal{B}=\{{\bf x}^{1},{\bf x}^{2},\ldots,{\bf x}^{n}\}$ be a minibatch of size $n$ with ${\bf x}^{i}=({\bf x}^{i}_{1},\ldots,{\bf x}^{i}_{d})$. The batch normalization transform is given by
@@ 3.8 Universal Approximation Theorem
Figure 3: Universal Approximation Theorem. A neural network with one hidden layer, three hidden units, and a $\tanh$ activation function approximates the function $f(x)=\sin(3x)$ in the interval $[-1,1]$. Fifty samples of the function $f(x)$ used to train the neural network are shown here as black dots. The orange line indicates the approximation, $\hat{f}(x)$, provided by the neural network.
So far, we have seen some fundamental aspects of deep learning, such as different types of neural networks and how to initialize them properly, the backprop algorithm, and batch normalization. Nevertheless, feed-forward neural networks are valuable tools in approximating functions. For instance, we can use a straightforward neural network of one hidden layer with three units to approximate function $f(x)=\sin(3x)$ in the interval $[-1,1]$. Figure 3 shows in black dots the data points we used to train the neural network,
@@
Neural networks with one hidden layer can, in theory, approximate any continuous function defined in a compact subspace of $\mathbb{R}$ per the Universal Approximation Theorem [17, 35]. More precisely, and according to [17], $I_{q}$ is the $q$-dimensional unit cube, $[0,1]^{q}$, and $C(I_{q})$ is the space of continuous functions on $I_{q}$, and $\alpha_{j},\beta_{j}\in\mathbb{R}$ are fixed parameters.
@@ 4 Solving Partial Differential Equations
For example, let the differential operator $\mathcal{T}$ be $-\frac{\partial^{2}}{\partial x^{2}}$ and take as boundary conditions $y(0,t)=y(1,t)=0$, and initial conditions $y(x,0)=y_{0}(x)$, then we can rewrite Equation (8) in one-dimension as:
@@
[…]
@@ 4.1 Deep Galerkin Method
$\begin{split}\mathcal{J}(\mathcal{F}({\bf x},t;\boldsymbol{\theta}))&=\underbrace{||(\partial_{t}+\mathcal{T})\mathcal{F}({\bf x},t;\boldsymbol{\theta})||_{[0,T]\times\Omega,\nu_{1}}^{2}}_{\text{Differential Operator}}+\underbrace{||\mathcal{F}({\bf x},t;\boldsymbol{\theta})-g({\bf x},t)||_{[0,T]\times\partial\Omega,\nu_{2}}^{2}}_{\text{Boundary Conditions}}+\underbrace{||\mathcal{F}({\bf x},0;\boldsymbol{\theta})-y_{0}({\bf x})||_{\Omega,\nu_{3}}^{2}}_{\text{Initial Conditions}},\end{split}$ (14)
@@ 31 L_boundary = ((y_bd1 - x_bd1)**2 + (y_bd2 - x_bd2)**2)
Figure 7: $1$D Heat equation solution. Panel A shows the exact solution of Equation (17), B illustrates the approximated solution by the DG method and a deep feed-forward neural network, and panel C displays the loss over training iterations.
@@ 4.3.1 Effects of Batch Size on Minimization
As we already have seen, a minimization problem, especially one involving neural networks, has many parameters that can affect the results. For instance, the size of the neural network, the number of hidden units, the learning rate, and many others are called hyperparameters to distinguish from the actual parameters of a neural network, which are the weights and the biases. We will discuss the hyperparameters later and see how to search for good sets of them using specialized software. For now, we will focus on one of those hyperparameters, the mini-batch size. We will investigate how the mini-batch size affects the minimization process and, thus, the loss function of our problem. To this end, we solve the problem of the one-dimensional heat Equation (17) using the loss function (4.3), and we follow the methodology we described in the previous section. Only this time, we vary the size of the mini-batches using powers of two, starting from $12^{0}=$ 1 and going up to $10242^{10}=$ 1024. We reduce the number of iterations to $1,000$, and we run for each mini-batch size the script that solves the one-dimensional heat equation. As we see in Figure 8, the slowest loss decrease occurs when the batch size is set to one. Once we start increasing the size of a mini-batch, the convergence of the minimization takes place faster. However, after a batch size of $4$, there is no practical convergence difference.
@@ 4.4 Search for the Hyperparameters
To use the Ray Tune, first, we have to write our objective function (the one we will maximize/minimize). In our case, we are looking for the hyperparameters that will minimize the loss function defined by Equation (4.3). The hyperparameters we aim to optimize are the batch size, the learning rate, and the number of iterations. Of course, we could search for the number of layers and units per layer along with the other three parameters. However, we would like to keep things as simple as possible. Thus, our objective function takes as argument a Python dictionary config that contains all the hyperparameters we are optimizing for. Then, we define a neural network (lines 33336666 in code listing 6), and we use that neural network to minimize our loss function (4.3). Finally, we transmit the last value of the loss to Ray Tune (line $16$) to be used in the hyperparameters optimization process. In this case, we are trying to find the set of hyperparameters that minimize the loss.
@@
In the code listing 7, we show how to use the objective function to search for the three hyperparameters: batch size (batch_size), number of iterations (n_iters), and learning rate (lrate). First, we define the search space and specify the distributions from which we will draw candidate values for our hyperparameters (lines 33336666). Then, we choose the search algorithm (line $9$), in this case, the Optuna [1] algorithm, and which scheduler we will use. Finally, we instantiate the Tuner class (lines 1414141426262626) and pass as arguments the resources we will use ($1$ GPU and $10$ CPU cores). We pass the objective function (objective) and the configuration, where we define the metric (loss in this case) and a minimization. Moreover, we pass the Optuna algorithm as an argument, the scheduler, and the number of samples for which the algorithm will perform the hyperparameters space search. Finally, we pass the search space (which is a Python dictionary), and we call the fit method. One can find more details on using Ray Tune on its documentation page [^7] https://docs.ray.io/en/latest/tune/index.html.

### v4

@@ 3.1 Feed-forward neural network
Computational units or neurons lie within each layer, performing affine transformations. We can write that in a vector form as ${\bf y}_{l}=f_{l}({\bf z})=g_{l}({\bf W}_{l}\cdot{\bf z}+{\bf b}_{l})$, where ${\bf W}_{l}$ contains the weights that connect neurons from layer l1l-1 with neurons in layer $l$ and ${\bf b}_{l}$ are the biases of the neurons at layer $l$, and number of layers $l$ runs $l=1,\ldots,k$. The function $g(x)$ is called the activation function and can be (i) a Sigmoid $g(x)=\frac{1}{1+\exp(-x)}$, (ii) a hyperbolic tangent $g(x)=\tanh(x)$, a (iii) rectified linear unit (ReLU) $g(x)=\max\{x,0\}$, any custom function, or any of the functions that the reader can find in Pytorch documentation page [^3] https://pytorch.org/docs/stable/nn.html##non-linear-activations-weighted-sum-nonlinearity.
@@
Like all the modern deep learning libraries, Pytorch offers all the necessary tools to build feed-forward neural networks and define affine transformations easily. The reader can find a complete list on the PyTorch documentation page [^4] https://pytorch.org/docs/stable/nn.html#linear-layers. Moreover, code listing 1 shows a basic implementation of a Pytorch Multi-Layer Perceptron neural network (a feed-forward neural network) with linear layers and $\tanh$ activation functions. We must define the layers and the activation functions used in the constructor __init__ method. The same method receives as arguments the input and output dimensions, which are the number of independent and dependent variables in our case, respectively. The argument input_dim defines the number of input units (independent variables such as $t$ and $\bf x$}), and on the other hand, the argument output_dim determines the number of units in the output layer (e.g., $y$). Moreover, it receives the number of layers (num_layers) and the number of units in each hidden layer (hidden_size). In this implementation, we assume that all hidden layers will have the same number of units. Keep in mind that if one has to use a neural network with one hidden layers, they have to set the number of layers to zero, for two hidden layers, the number has to be set to one, and so on. In line $11$ of code listing 1, we define the input layer; in lines 14141515, the hidden layers; and in line $18$, the output layer. The class Linear of Pytorch is an implementation of an affine transformation (${\bf y}={\bf W}\cdot{\bf x}+{\bf b}$). Notice that in line $14$, we use a ModuleList to instantiate the hidden layers. We do so because the number of hidden layers is not known a priori; instead, it is passed as an argument to the constructor when we instantiate the MLP class. The forward method takes the input tensor as an argument (it might accept more arguments, as we will see later) and applies all the transformations and activation functions we have already defined. Essentially, this method runs the forward pass of our neural network, which is the part of the code where we determine the structure of our neural network. Keep in mind that when we use the batch normalization (batch_norm = True), we can eliminate the bias term from the previous Layer by setting bias=False in the constructor of the Layer class in Pytorch.
1class MLP($n$.Module):
@@
11 self.fc_in = $n$.Linear(input_dim, hidden_size)
@@
14 self.layers = $n$.ModuleList([$n$.Linear(hidden_size, hidden_size)
@@
18 self.fc_out = $n$.Linear(hidden_size, output_dim)
@@
25 self.bn = $n$.BatchNorm1d(hidden_size)
@@ 3.2 DGM neural network
1 class DGMLayer($n$.Module):
@@
5 self.Z_wg = $n$.Linear(hidden_size, hidden_size)
6 self.Z_ug = $n$.Linear(input_dim, hidden_size, bias=False)
@@
8 self.G_wz = $n$.Linear(hidden_size, hidden_size)
9 self.G_uz = $n$.Linear(input_dim, hidden_size, bias=False)
@@
11 self.R_wr = $n$.Linear(hidden_size, hidden_size)
12 self.R_ur = $n$.Linear(input_dim, hidden_size, bias=False)
@@
14 self.H_wh = $n$.Linear(hidden_size, hidden_size)
15 self.H_uh = $n$.Linear(input_dim, hidden_size, bias=False)
@@
18 self.sigma = $n$.Tanh()
@@ 3.3 Feed-forward ResNet
An example of implementing a linear ResNet block in Pytorch is given in code listing 3. The method __init__ takes the number of units in the input and output layers as arguments, respectively. The argument downsample defines a transformation (i.e., ${\bf W}_{s}$) at the skip connections that we will apply when the input’s $\bf x$} dimensions mismatches the size of the output of the residual block (see Figure 2 C). Finally, the argument num_features indicates if the input tensor will contain an extra dimension besides the batch size and the independent variables. Notice that the linear layers here are defined using the Pytorch Sequential container in lines 10101414. Within each container, we instantiate a Pytorch linear layer followed by a batch normalization layer (see Section 3.7). In the forward method in lines 23233333, we apply two linear layers (see lines $20$ and $21$ in code listing 3) and linear rectifications on their outputs. Then we check if downsample is not None (see lines $23$ and $24$), and if it’s not, then we apply the downsample transformation to the input $\bf x$}, otherwise our residual remains the original input $\bf x$}. Once we have the residual values, we add them to the output, apply linear rectification, and return the output.
1class ResidualBlock($n$.Module):
@@
15 self.fc1 = $n$.Sequential($n$.Linear(input_dim, output_dim, bias=False),
16 $n$.BatchNorm1d(self.num_features))
@@
18 self.fc2 = $n$.Sequential($n$.Linear(output_dim, output_dim, bias=False),
19 $n$.BatchNorm1d(self.num_features))
@@
21 self.relu = $n$.ReLU()
@@ 3.5 Backpropagation, Optimizers and Automatic Differentiation
Algorithm 1 summarizes the learning steps, and code listing 4 shows the backbone of training a neural network in Pytorch. In general, we first instantiate our neural network class (line $2$ in code listing 4), then we specify the optimizer (in this case, at line $4$ we use the Adam optimizer [40]), and the criterion, which is the loss function we are interested in minimizing. In this work, we define custom loss functions per equation; we will see in Section 4.1 how to implement a custom loss function. Moreover, we might want to initialize a learning rate scheduler, which will change (usually decrease) the learning rate based on some linear, exponential, sinusoidal, or other relation [^6] The reader can find a list of learning rate schedulers already implemented in PyTorch here: https://pytorch.org/docs/stable/optim.html#how-to-adjust-learning-rate. Then, for each iteration (line $7$), we sample our independent variables (e.g., time $t$, space $x$, or both) based on predefined distributions. Once we get the samples, we zero the grads (line $10$), then we approximate the solution yhat using the neural network and the independent variables. Finally, we pass the approximation to the loss function, do a backward step (lines 1313-$18$), and finally minimize it (line $21$).
@@ 3.7 Batch Normalization
Let $\mathcal{B}=\{{\bf x}^{1},{\bf x}^{2},\ldots,{\bf x}^{n}\}$ be a minibatch of size $n$ with ${\bf x}^{i}=({\bf x}^{i}_{1},\ldots,{\bf x}^{i}_{d})$. The batch normalization transform is given by
@@ 3.8 Universal Approximation Theorem
Figure 3: Universal Approximation Theorem. A neural network with one hidden layer, three hidden units, and a $\tanh$ activation function approximates the function $f(x)=\sin(3x)$ in the interval [1,1][-1,1]. Fifty samples of the function $f(x)$ used to train the neural network are shown here as black dots. The orange line indicates the approximation, $\hat{f}(x)$, provided by the neural network.
So far, we have seen some fundamental aspects of deep learning, such as different types of neural networks and how to initialize them properly, the backprop algorithm, and batch normalization. Nevertheless, feed-forward neural networks are valuable tools in approximating functions. For instance, we can use a straightforward neural network of one hidden layer with three units to approximate function $f(x)=\sin(3x)$ in the interval [1,1][-1,1]. Figure 3 shows in black dots the data points we used to train the neural network,
@@
Neural networks with one hidden layer can, in theory, approximate any continuous function defined in a compact subspace of $\mathbb{R}$ per the Universal Approximation Theorem [17, 35]. More precisely, and according to [17], $I_{q}$ is the $q$-dimensional unit cube, $0,1]^{q}$, and $C(I_{q})$ is the space of continuous functions on $I_{q}$, and $\alpha_{j},\beta_{j}\in\mathbb{R}$ are fixed parameters.
@@ 4 Solving Partial Differential Equations
For example, let the differential operator $\mathcal{T}$ be $\frac{\partial^{2}}{\partial x^{2}}$ and take as boundary conditions $y(0,t)=y(1,t)=0$, and initial conditions $y(x,0)=y_{0}(x)$, then we can rewrite Equation (8) in one-dimension as:
@@
$-\int_{\Omega}\nabla v\cdot\nabla\hat{y}\,dV$ $=\int_{\Omega}vfdV,\quad\forall v\in\mathcal{X}_{0}^{K},$ (13)
@@ 4.1 Deep Galerkin Method
where the first term tells us how good our approximation of the differential operator is, and the second and third terms indicate how well we approximate the boundary and the initial values, respectively. The norm is an $L^{2}$ given by $f(y)||_{\mathcal{Y},\nu}^{2}=\int_{\mathcal{Y}}|f(y)|^{2}\nu(y)dy$ with $\nu$ being a density defined on $\mathcal{Y}$. In Equation (15), $\mathcal{Y}$ becomes $0,T]\times\Omega$, $0,T]\times\partial\Omega$, and $\Omega$ in the three terms of the right-hand side, respectively. By minimizing Equation (15), we obtain an approximated solution to the IVBP described by Equation (8) (i.e., when $\mathcal{J}(\mathcal{F}(\boldsymbol{\theta};{\bf x},t))=0$, then the solution $\hat{y}$ provided by the neural network $\mathcal{F}(\boldsymbol{\theta};{\bf x},t)$ will satisfy the IVBP (8)). Therefore, our main objective is to find a set of parameters $\boldsymbol{\theta}$ of a neural network $\mathcal{F}(\boldsymbol{\theta};{\bf x},t))$ that will minimize the loss function described by Equation (15). However, and accordinly to [60], this is not always possible since when the dimension $q$ ($q$ is the dimension of domain and variables, see Section 2) of the problem increases, the integrals over $\Omega$ become intractable. To overcome this problem we employ a machine learning method where we can randomly draw samples from distributions on $\Omega$ and $\partial\Omega$ and estimate the squared error:
@@ 31 L_boundary = ((y_bd1 - x_bd1)**2 + (y_bd2 - x_bd2)**2)
Figure 7: 11D Heat equation solution. Panel A shows the exact solution of Equation (17), B illustrates the approximated solution by the DG method and a deep feed-forward neural network, and panel C displays the loss over training iterations.
@@ 4.3.1 Effects of Batch Size on Minimization
As we already have seen, a minimization problem, especially one involving neural networks, has many parameters that can affect the results. For instance, the size of the neural network, the number of hidden units, the learning rate, and many others are called hyperparameters to distinguish from the actual parameters of a neural network, which are the weights and the biases. We will discuss the hyperparameters later and see how to search for good sets of them using specialized software. For now, we will focus on one of those hyperparameters, the mini-batch size. We will investigate how the mini-batch size affects the minimization process and, thus, the loss function of our problem. To this end, we solve the problem of the one-dimensional heat Equation (17) using the loss function (4.3), and we follow the methodology we described in the previous section. Only this time, we vary the size of the mini-batches using powers of two, starting from $2^{0}=1$ and going up to $2^{10}=1024$. We reduce the number of iterations to $1,000$, and we run for each mini-batch size the script that solves the one-dimensional heat equation. As we see in Figure 8, the slowest loss decrease occurs when the batch size is set to one. Once we start increasing the size of a mini-batch, the convergence of the minimization takes place faster. However, after a batch size of $4$, there is no practical convergence difference.
@@ 4.4 Search for the Hyperparameters
To use the Ray Tune, first, we have to write our objective function (the one we will maximize/minimize). In our case, we are looking for the hyperparameters that will minimize the loss function defined by Equation (4.3). The hyperparameters we aim to optimize are the batch size, the learning rate, and the number of iterations. Of course, we could search for the number of layers and units per layer along with the other three parameters. However, we would like to keep things as simple as possible. Thus, our objective function takes as argument a Python dictionary config that contains all the hyperparameters we are optimizing for. Then, we define a neural network (lines 3366 in code listing 6), and we use that neural network to minimize our loss function (4.3). Finally, we transmit the last value of the loss to Ray Tune (line $16$) to be used in the hyperparameters optimization process. In this case, we are trying to find the set of hyperparameters that minimize the loss.
@@
In the code listing 7, we show how to use the objective function to search for the three hyperparameters: batch size (batch_size), number of iterations (n_iters), and learning rate (lrate). First, we define the search space and specify the distributions from which we will draw candidate values for our hyperparameters (lines 3366). Then, we choose the search algorithm (line $9$), in this case, the Optuna [1] algorithm, and which scheduler we will use. Finally, we instantiate the Tuner class (lines 14142626) and pass as arguments the resources we will use ($1$ GPU and $10$ CPU cores). We pass the objective function (objective) and the configuration, where we define the metric (loss in this case) and a minimization. Moreover, we pass the Optuna algorithm as an argument, the scheduler, and the number of samples for which the algorithm will perform the hyperparameters space search. Finally, we pass the search space (which is a Python dictionary), and we call the fit method. One can find more details on using Ray Tune on its documentation page [^7] https://docs.ray.io/en/latest/tune/index.html.